OpenStudy (anonymous):

can somone help me again please i have 15 3/5 + 9 1/2 my answer was 24 51/10 then simpliest form was 29 1/10 but my answer key says 25 1/10 can someone tell me where i went wrong??

OpenStudy (anonymous):

why don't you first convert the mixed fraction --> into a fraction then find the same denominators and solve? ^_^

OpenStudy (anonymous):

when you do that, you'll get the following form:\[15 \frac{3}{5} + 9 \frac{1}{2} = \frac{78}{5} + \frac{19}{2}\] find the same denominators and use the concept that I have showed you in the previous question :) give it a try now ^_^

OpenStudy (anonymous):

ok so i got 10 as the common den. then did 78*2=156 and 19*5=95 so that gives me 156/10 +95/10=251/10. Then i divided 251 by 10 and got 25 with a remainder of 1 and then i came up with 25 1/10 right??

OpenStudy (anonymous):

so the remainder is always the top number and the one i divided by is the denominator and the answer is the whole number??

OpenStudy (anonymous):

excellent!

OpenStudy (anonymous):

^_^ lol, all you've got to do is give yourself a chance, you're smart

OpenStudy (anonymous):

k and it works the same way with subtraction??

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

basic idea, you have to have same denominators ~ that's all ^_^

OpenStudy (anonymous):

Thank u again so much i am takin my entance exam to nursin school today at 5 and fractions r the hardest for me just tryin to work out the kinks u have been suck a HUGE help to me!!

OpenStudy (anonymous):

then just calculate normally

OpenStudy (anonymous):

a pleasure dear :) anytime ^_^

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

THANKS!!